OpenStudy (anonymous):

The sum of two numbers is 48. If one third of one number is 5 greater than one sixth of another number, which of the following is the smaller number? Help me please...

OpenStudy (decarr432):

Welcome to openstudy!

OpenStudy (decarr432):

Are there any answer choices given?

OpenStudy (mathstudent55):

Do you need to write a system of equations and then solve it?

OpenStudy (radar):

Use algebra. Label your variable Let x = value of one number then, 48-x is the value of the other number. This is true because their sum is 48 and was given. Now write their equivalent relationships into an algebraic equation. x/3 + 5 = (48 - x)/6 Now solve for x ......it is that simple.

OpenStudy (radar):

Now I rewrite the equation: x/3 + 5 = 8 - x/6 all I did here was the the quantity in the parentheses by the 6 I now subtract 5 from both sides of the equal sign, getting: x/3 = 8 -5 - x/6, I now add x/6 to both sides. Getting: x/3 + x/6 = 8 - 5 Now simplify by doing the required arithmetic operations on the equation: 1/2 x = 3 Multiplying both sides by 2 to get 1x x = 6 one of the values and the smallest the other value would be 48 - x or 48 - 6 = 42 Good luck with your studies.

OpenStudy (mathstudent55):

@radar Your second equation is incorrect. Check your numbers (6 and 42) with the original statements of the problem, and you'll see that they don't work. This is your mistake: In general, if a is 5 greater than b, then a = b + 5. You did a + 5 = b. If 1/3 of x is 5 greater than 1/6 of y, then 1/3 x = 1/6 y + 5 You have 1/3 x + 5 = 1/6 y

OpenStudy (mathstudent55):

You can do this also with a system of equations. Let x = one number and let y = the other number You can get an equation out of this statement. The sum of two numbers is 48. \(x + y = 48\) You get the second equation out of this statement. If one third of one number is 5 greater than one sixth of another number \(\dfrac{1}{3}x = \dfrac{1}{6}y + 5\) Now we have a system of equations: \(x + y = 48\) \(\dfrac{1}{3}x = \dfrac{1}{6}y + 5\) We multiply both sides of the second equation by 6 to get rid of the denominators. \(x + y = 48\) \(6 \times \dfrac{1}{3}x = 6 \times \dfrac{1}{6}y + 6 \times 5\) \(x + y = 48\) \(2x = y + 30\) Subtract y from both sides of the second equation. \(x + y = 48\) \(2x -y = 30\) Add the equations to eliminate y. 3x = 78 x = 26 Substitute x = 26 in the first original equation and solve for y. x + y = 48 26 + y = 48 y = 22 We have one number is 26 and the other number is 22. Now let's check to see if our solution makes sense. "The sum of two numbers is 48." \(22 + 26 = 48\) Yes, the sum of the numbers is 48. That makes the first statement true. Now let's check the second statement. "one third of one number is 5 greater than one sixth of another number" \(\dfrac{26}{3} = \dfrac{22}{6} + 5\) \(\dfrac{26}{3} = \dfrac{11}{3} + \dfrac{15}{3} \) \(\dfrac{26}{3} = \dfrac{26}{3} \) This makes the second statement true. The numbers are 26 and 22.

OpenStudy (radar):

@mathstudent55 Thanks, you know I thought I was off on that, but went ahead any way. I did check my answer and with my misinterpretation of which was greater verified what I did, the only problem is....I did the wrong thing lol.

OpenStudy (radar):

@mathstudent By the way, your explanation with the step by step was great

OpenStudy (radar):

@sawyer345 Sorry if I erred (which I did), I hope you saw mathstudent55 solution and now understand.

OpenStudy (mathstudent55):

@radar Thanks for your kind words.